OpenStudy (nastech):

prove that sin3A=3sinA-(4sin^3)A

OpenStudy (campbell_st):

use the sum of 2 angles expansion sin(A + 2A) = sin(A)cos(2A) + cos(A)sin(2A) now sin 2A = 2sin(A)cos(A) cos(2A) = 1 - 2sin^2(A) make the substitutions and it should fall out